OpenStudy (anonymous):

for the function f, construct and simplify the difference quiotent f(x+h)-f(x) ------ h f(x)=x/4-x

OpenStudy (marcelie):

omgg im learning this... what college do you go too?

OpenStudy (anonymous):

\[\frac{ f(x+h)-f(x) }{ h }\]

OpenStudy (whpalmer4):

I assume that is \[f(x)=\frac{x}{4-x}\]

OpenStudy (anonymous):

valencia, and yes that is

OpenStudy (whpalmer4):

ok, so replace \(x\) with \(x+h\) in \(f(x)\), what do you get?

OpenStudy (anonymous):

\[\frac{ x+h }{ 4-x+h }\]

OpenStudy (whpalmer4):

not quite. \[4-(x+h)=\]

OpenStudy (anonymous):

\[4(-x-h)\]

OpenStudy (whpalmer4):

\[4-(x+h)=4-x-h\]right?

OpenStudy (anonymous):

yes!

OpenStudy (whpalmer4):

ok, so \[f(x+h)=\frac{x+h}{4-x-h}\] \[f(x)=\frac{x}{4-x}\] \[\frac{f(x+h)-f(x)}{h}=\]

OpenStudy (whpalmer4):

some tedious but straightforward algebra to do...

OpenStudy (anonymous):

so i need to get rid of the "h" by multiplying ?

OpenStudy (whpalmer4):

no, you need to make a common denominator so you can add the two fractions, then divide the mess by h

zepdrix (zepdrix):

Valencia? Ooo nice :) I'm over at UCF.